Can you please help with this equation: Write complex number in polar form: √7+√21 • i
Answer by stanbon(75887)

Write complex number in polar form: √7+√21 • i
Complex Number form:: x + yi
----
Your Problem::
x = sqrt(7)
y = sqrt(21)
------
Converting to Polar:
r = sqrt[x^2+y^2] = sqrt[7+21] = sqrt(28) = 2sqrt(7)
-----------
theta = arctan(y/x) = arctan(sqrt(21)/sqrt(7) = arctan(sqrt(3)) = pi/3
-----------------------
Ans: (2sqrt(7),pi/3)
--------------------------
Cheers,
Stan H